# Geometric decomposition of planar vector fields with a limit cycle

### v1

@@ 1 Introduction
Potential systems often present single point attractors, known as fixed points, to which the system evolves (see Figure 1). However, more complicated behaviours cannot be observed in such systems and this included notably oscillatory behaviour, as the system eventually settles at fixed points, or to infinity [9]. studied Morse-Smale potential systems and they did not pursue the analysis of phenomena such as periodic oscillations.
@@
Periodic trajectories in Hamiltonian systems are neither unstable nor stable and perturbations will induce the system to follow a different orbit, in the same way that a push of a swing will increase or reduce the amplitude. These are in stark contrast with biological oscillations, which exhibits (stable) limit cycle behaviour: the system evolves to a single orbit of constant size and perturbations always return to such orbit (see Figure 1A) [9]. However, neither Hamiltonian nor potential systems, at least without time dependence, can fully describe limit cycle behaviour [9, 10].
@@
where $\xi(\bm{x},t)$ is a noise term. Under certain assumptions, one can compute the potential, as well as both matrices. Although some authors have claimed this could be seen as the HHD when there is no noise, this might not be always true [14, 16].
There have been other attempts at decomposing limit cycle dynamics into potential and rotational elements, these methods rely on using stochastic properties, such as diffusion [17, 18]. Nonetheless, this line of research is still contained within the field of SDEs, and justifications for the different assumptions are often made based on statistical arguments [15, 19]. Still, the existence of such decompositions or a latent geometric structure should be feasible also in Ordinary Differential Equations (ODE), and should reveal the most fundamental properties of deterministic models.
@@
To this day, the application of geometric decompositions to study limit cycle oscillators in biology is still in its infancy despite the large body of research in related problems and the potential to deepen our understanding of biological systems [23]. Additionally, the computation of solutions for these formalisms often restricted to analytical expressions of idealised systems.
@@
The main objective of this work is finding a reformulation of such a system in geometric terms, giving a more useful and insightful representation. In other words, we seek to find a different description of a system with two variables (mRNA, proteins, species) to a new set of variables (phase and amplitude in the case of biological oscillators) that facilitates comparison and analysis. I operate under the assumption that the problem admits such a formulation, despite not proving it.
Mathematical notation will be as compact as possible and respecting conventions in different fields when possible. To ensure this paper can reach a broader audience, it might not as be rigorous as desired for certain readers while being too mathematical for others. From a mathematical perspective, I will assume the level of Strogatz [9]. A general overview on vector calculus can be found on Appendix A.
The paper is structured as follows: In Section 2, I will derive the decomposition as well as give some geometric interpretation. In Section 3, I will give further insight into the formalism using a relatively simple nonlinear model. In Section 4, I will apply the present decomposition to linear(ised) systems in an analytic way. Finally, in Section 5, I will present an algorithm to compute the decomposition numerically for any system with a single limit cycle enclosing a fixed point and apply to a particular case of interest.
@@ 4 Analytical decomposition of a linear system
Before moving to the general case, and to fully understand the geometric meaning of Equation 6, I will apply it to a linear system or, more generally, to a linearised system around a fixed point. Without loss of generality, I will assume such fixed point is centered at the origin and it is described as
@@
However, this decomposition is not unique: $J_{0}$ remains the same if $\kappa_{0}$ is divided by a factor $k$ and $F_{0}$ is multiplied by the same factor. However, uniqueness can be achieved by defining
@@ 6 Discussion
Although a direct method could be done by solving a Hamilton-Jacobi type equation, a large number of challenges emerge in this type of problems, from unknown boundary conditions to numerical instability. However, the present method is a simple alternative to compute the proposed decomposition by transforming the problem into a collection of limit cycles. This process yields sufficiently good results, although further research is required. Furthermore, the method could be modified such that integration is no longer required.

### v2

@@ 1 Introduction
Potential systems often present single point attractors, known as fixed points, to which the system evolves (see Figure 1). However, more complicated behaviours cannot be observed in such systems and this included notably oscillatory behaviour, as the system eventually settles at fixed points, or to infinity [9]. Rand and coworkers studied Morse-Smale potential systems and they did not pursue the analysis of phenomena such as periodic oscillations [2].
@@
Periodic trajectories in Hamiltonian systems are neither unstable nor stable and perturbations will induce the system to follow a different orbit, in the same way that a push of a swing will increase or reduce the amplitude. These are in stark contrast with biological oscillations, which exhibits (stable) limit cycle behaviour: the system evolves to a single orbit of constant size and perturbations always return to such orbit (see Figure 1A) [9]. However, neither Hamiltonian nor potential systems, at least without time dependence or adding new dummy variables, can fully describe limit cycle behaviour [9, 10].
@@
where $\xi(\bm{x},t)$ is a noise term. Under certain assumptions, one can compute the potential, as well as both matrices. However, it is worth noting that the computation of these matrices requires the definition of a diffusion matrix, based on the structure of the noise. Although some authors have claimed this could be seen as the HHD when there is no noise, this might not be always true [14, 16].
However, the SA-SDE decomposition can be seen as a reinterpretation of the Freidlin-Wentzell (FW) Potential for SDEs [17, 5, 18]. In this framework, the FW potential relates to the probability of a trajectory escaping a certain basin of attraction.
There have been other attempts at decomposing limit cycle dynamics into potential and rotational elements, these methods rely on using stochastic properties, such as diffusion [19, 20]. Nonetheless, this line of research is still contained within the field of SDEs, and justifications for the different assumptions are often made based on statistical arguments [15, 21, 18]. Still, the existence of such decompositions or a latent geometric structure should be feasible also in Ordinary Differential Equations (ODE), and should reveal the most fundamental properties of deterministic models. In other words, it should be possible to describe a deterministic dynamical system without invoking any kind of probabilistic argument or stochastic properties. Furthermore, one could analyse the system using the tools and concepts of differential geometry, such as metric tensors or differential forms [22].
@@
To this day, the application of geometric decompositions to study limit cycle oscillators in biology is still in its infancy despite the large body of research on related problems and the potential to deepen our understanding of biological systems [26]. Additionally, the computation of solutions for these formalisms often restricted to analytical expressions of idealised systems.
@@
The main objective of this work is finding a reformulation of such a system in geometric terms, giving a more useful and insightful representation. In other words, we seek to find a different description of a system with two variables (mRNA, proteins, species) to a new set of variables (phase and amplitude in the case of biological oscillators) that facilitates comparison and analysis. I operate under the assumption that the problem admits such a formulation, despite not proving it. However, similar work of FW potentials can be seen as the rigorous formulation of the problem [18].
The paper is structured as follows: In Section 2, I will derive the decomposition as well as give some geometric interpretation. In Section 3, I will give further insight into the formalism using a relatively simple nonlinear model. In Section 4, I will apply the present decomposition to linear(ised) systems in an analytic way. Finally, in Section 5, I will present a novel algorithm to compute the decomposition numerically for any system with a single limit cycle enclosing a fixed point and apply to a particular case of interest.
@@ 4 Analytical decomposition of a linear system
Before moving to the general case, and to fully understand the geometric meaning of Equation 6, I will apply it to a linear system or, more generally, to a linearised system around a fixed point. The following results can be seen as a reformulation of those in [18], although from a purely deterministic perspective. However, final results might differ as a consequence of different constrains.
Without loss of generality, I will assume such fixed point is centered at the origin and it is described as
@@
However, this decomposition is not unique: $J_{0}$ remains the same if $\kappa_{0}$ is divided by a factor $k$ and $F_{0}$ is multiplied by the same factor. This is a consequence of the covariance between both objects and it can serve as the basis of an approach based on differential geometry, in which $\kappa_{0}$ plays the role of a metric tensor [22, 30, 31]
If uniqueness is a property to be desired, then two approaches can be taken. The first one is setting the diagonal terms of the normal matrix to one and then proceed to compute the corresponding gradient term, which corresponds to the FW potential [18]. However, this approach can lead to singularities for Hamiltonian systems, where $\kappa_{0}$ becomes undefined. Therefore, an alternative definition for this matrix is
@@ 6 Discussion
Alternatively, the lack of uniqueness can be seen as a consequence of the general covariance between the product of a metric tensor and a gradient [22]. Under this framework, the determination of a unique solution based on constrains is just the choice of a gauge. This built-in flexibility should be seen as a feature (and possibly beneficial) and thus the uniqueness would emerge as a consequence of enforcing a perspective on the system. In this way, both the FW and the SA-SDE potentials are nothing but two valid descriptions of the same system, given complementary explanations of the system.
Interestingly, the geometry described by this metric is no longer Riemannian nor Symplectic, but Finsler [30, 31]. In this case, the Finsler metric is basically a Riemannian metric without the quadratic restriction [30]. This opens interesting new ways to study dynamical systems, in particular biological ones, as Riemannian Geometry is still the dominant framework [2, 1].
@@
Although a direct method could be done by solving a Hamilton-Jacobi type equation, a large number of challenges emerge in this type of problems, from unknown boundary conditions to numerical instability. Several methods have been proposed to solve these problems [18, 34, 35, 33].
The present method is a simple alternative to compute the proposed decomposition by transforming the problem into a collection of limit cycles. Although this method might not be better than other methods in the literature, this is a novel approach, which could be serve as the basis of better algorithms. This process yields sufficiently good results, although further research is required. Furthermore, the method could be modified such that integration is no longer required.